OpenStudy (anonymous):

how many grams of chlorine are needed to make 56.6 g of aluminum chloride

OpenStudy (abmon98):

2Al+3Cl2-->2AlCL3 number of moles=mass/Molar mass number of moles=56.6/(27.0+(35.5*3)=0.42 moles Ratio of Cl2 to AlCl3 is 3:2 0.42*3/2=0.635 0.635*71=45.2 grams